A cylindrical copper wire with radius a = 0.021-m carries a uniform current of 229-A. What is the magnitude of the magnetic field at r = 0.0-m, r = a/2 and r = 2·a?
Answer #1

magnitude of the magnetic field at r = 0.0-m   = 0 T

=> magnitude of the magnetic field at r = a/2   =   2 * 10-7 * 229 * (0.25)/0.0105

                                                                         = 1.09 * 10-3 T

=> magnitude of the magnetic field at r = 2a =   2 * 10-7 * 229 /0.042

                                                                         = 1.09 * 10-3 T